OpenStudy (anonymous):

Solve |6x – 5| = 25.

OpenStudy (anonymous):

x=5, -10/3

OpenStudy (zarkon):

\[6x-5=25\] \[-(6x-5)=25\]

OpenStudy (zarkon):

solve both

OpenStudy (anonymous):

6x-5=25 x = 30/6=5 6x-5=-25 x = -20/6=-10/3